John is 13 years less than 3 times as old as his son. Six years ago, he was 14 years more than twice as old as his son was then. How old is John?
Answer by stanbon(75887)

John is 13 years less than 3 times as old as his son.
Six years ago, he was 14 years more than twice as old as his son was then.
How old is John?
-------------------------
Equations:
j = 3s-13
(j-6) = 2(s-6)+14
---------------------
Substitute for "j" and solve for "s":
(3s-13-6) = 2s-12+14
3s-19 = 2s+2
s = 21 (son's age now)
---
Since j = 3s-13, j = 3(21)-13 = 50 (john's age now)
===================================================
Cheers,
Stan H.
===================================================